OpenStudy (anonymous):

1.Solve S = 4v2 for v.

OpenStudy (lgbasallote):

divide both sides by 4 to get rid of the coefficient of c \[\frac{S}{4} = v^2\] now take the square root to get just v ;)

OpenStudy (lgbasallote):

^that's supposed to be coefficient of v*

OpenStudy (amistre64):

if you could tell us where you are having troubles with this, it would be helpful.

OpenStudy (anonymous):

I"m Confused.

OpenStudy (amistre64):

on which part, divide or square root?

OpenStudy (anonymous):

divide

OpenStudy (amistre64):

since "v" is stuck inside alot of mathing operations; we have to undo those operation to get to "v" by itself division undoes multiplication.

OpenStudy (amistre64):

for example: 5*4 = 20 5*4/4 = 5